OpenStudy (anonymous):

what is the number of solutions to the system? x + y = -1 5x - 5y = 7

OpenStudy (anonymous):

someone please help me! i dont get this

OpenStudy (m):

solve for x then plug it in the second equation

OpenStudy (anonymous):

so i would set it up like x + y = -1 = 5x - 5y = 7 ?

OpenStudy (m):

x = -1 - y 5(-1 - y) -5y = 7 then you'll know what y is

OpenStudy (m):

if you know y then you can plug in y to first equation and find out what x is

OpenStudy (anonymous):

wait let me write this down and try to do that

OpenStudy (m):

-5 - 5y -5y = 7 -10y = 14 y = -14/10

OpenStudy (m):

when you have two unknown (here x and y) then you always want to solve for x or y, then plug it into the other equation so you're working with just one unknown variable

OpenStudy (anonymous):

ok i get the -10 y part. buy what about the -5 and where did 14 come from?

OpenStudy (anonymous):

ahhh. isee.

OpenStudy (m):

you try to get y on left and everything else on right

OpenStudy (anonymous):

yeah but you left -5 out. what did you do with it? and where did 14 come from?

OpenStudy (m):

you move -5 to right side of = sign

OpenStudy (m):

you want to get rid of -5 on left side so you add +5 to both side

OpenStudy (m):

so -5 disappears on left and +5 appears on right

OpenStudy (anonymous):

Okay. but that equals 12.

OpenStudy (m):

whatever you do on left you MUST do same on right to keep equation same

OpenStudy (m):

sorry it's 12 :)

OpenStudy (m):

12/10

OpenStudy (anonymous):

i know that. ahh seee. thaats why i was like huh? lol

OpenStudy (anonymous):

u got me confused rii there. its cool tho

OpenStudy (m):

so plug in y= -12/10 to find x

OpenStudy (anonymous):

isnt it 12/-10 ?

OpenStudy (m):

it doesn't matter if it's on top or bottom

OpenStudy (m):

it's (-1) (12/10)

OpenStudy (m):

^^ that's the correct form

OpenStudy (anonymous):

ahhhhh i seee. so the answer is one solution.

OpenStudy (m):

there's only one solution for y and one for x

OpenStudy (m):

answer is a fraction

OpenStudy (anonymous):

i know. but its asking me how many solutions. one two or no solutions

OpenStudy (anonymous):

i have this other one. i have to find the ordered pairs. let me see if i can do it. -3x + 5y = 4 5x +5y = 60

OpenStudy (m):

i'm not sure what the question is asking?

OpenStudy (m):

there can be infinite combination of x and y and it'll give you those answers

OpenStudy (anonymous):

it says give the ordered pairs that satisfies the system of equations

OpenStudy (anonymous):

well for the other one that your talking about.... it was asking WHAT IS THE NUMBER OF SOLUTIONS TO THE SYSTEM? and i put ONE SOLUTION

OpenStudy (anonymous):

theres another choice that says INFINITELY MANY SOLUTIONS

OpenStudy (m):

i think it may be asking for just x and y like (x, y)

OpenStudy (m):

so just one solution

OpenStudy (anonymous):

yeah it is.

OpenStudy (anonymous):

Okay. thanks.

OpenStudy (anonymous):

let me see if i can do the other one.

OpenStudy (anonymous):

i have a question

OpenStudy (anonymous):

i did 5x ( -3x + 5y ) +5y = 60

OpenStudy (m):

no you have to solve for y first

OpenStudy (m):

y = (1/5)(3x +4)

OpenStudy (m):

then you plug it in like you did

OpenStudy (anonymous):

so thats how im suppose set it up rii?

OpenStudy (m):

i'm not sure what you did

OpenStudy (m):

you first either solve for x or y

OpenStudy (m):

then plug in the result to the second equation

OpenStudy (anonymous):

i want to solve for X

OpenStudy (anonymous):

so how do i set it up?

OpenStudy (m):

x = (5y - 4)(1/3)

OpenStudy (m):

5[(5y - 4)(1/3)] +5y = 60 the bracket [ ] is the x

OpenStudy (m):

now you only have one variable y in your equation since we plugged in for x

OpenStudy (anonymous):

where did one third come from?

OpenStudy (anonymous):

wait let me get this straightt....

OpenStudy (m):

i just plugged in for x

OpenStudy (anonymous):

ahh okay

OpenStudy (anonymous):

so now that we have our equation we have to distribute the 5 to the perenthesis (5y-4) correct?

OpenStudy (m):

you can do that then divide everything by 3

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (anonymous):

i can distribute 5 to 1/3 righT?

OpenStudy (m):

you're not distributing since (1/3) is its own. you simply multiply

OpenStudy (anonymous):

srry about that

OpenStudy (anonymous):

so what i shud have is [ (25y-20) (1/3) ]

OpenStudy (anonymous):

[ (25y-20) (1/3) ] + 5y = 60

OpenStudy (m):

yeah then solve for y

OpenStudy (m):

now you have to distribute (1/3)

OpenStudy (anonymous):

im a little confused. so right this is what i have...... [ (25y-20) (1/3) ] + 5y = 60 -5y - 5y [ (20y- 20) (1/3) ] = 60

OpenStudy (m):

no

OpenStudy (m):

you don't want to move 5y to other side because you still have 25y

OpenStudy (m):

you want to move 20(1/3) to right

OpenStudy (m):

then combine 25y(1/3) + 5y

OpenStudy (m):

you may want to use calculator

OpenStudy (anonymous):

so multiply -20/2 times 1/3

OpenStudy (m):

remember we're trying to solve for y

OpenStudy (anonymous):

wait ignore that.. typo

OpenStudy (anonymous):

so multiply 20/1 times 1/3 then combien 25y times 1/3 + 5y then combine like terms so you'd end up with 30y (1/3)

OpenStudy (m):

no

OpenStudy (m):

you first distribute (1/3) to 25y and -20

OpenStudy (m):

so it's 25y(1/3) + 5y = 60 + 20(1/3)

OpenStudy (m):

y(25(1/3) + 5) = 60 + 20(1/3) y = 60 + 20(1/3) / (25(1/3) + 5)

OpenStudy (m):

it looks real messy b/c i didn't simplify anything

OpenStudy (m):

\[y = 60 + 20(1/3)\div (25(1/3) + 5)\]

OpenStudy (anonymous):

so now im gunna distribute 20 (1/3) and 25 (1/3)

OpenStudy (m):

no you just multiply

OpenStudy (m):

it's just 20/3 and 25/3

OpenStudy (m):

you can multiply by (1/3) or just divide by 3

OpenStudy (anonymous):

500\3

OpenStudy (m):

where'd that come from